B) (b) Use technology to compute the standard error, the point estimate for σ.
se = ??? (Round to four decimal places as needed.)
(c) Assuming the residuals are normally distributed, use technology to determine sb . 1
sb1= ??? (Round to four decimal places as needed.)
(d) Assuming the residuals are normally distributed, test H0: β1 = 0 versus H1: β1 ≠ 0 α = 0.05 level of significance.
Use the P-value approach.
Determine the P-value for this hypothesis test.
P-value = ??? (Round to three decimal places as needed.)For the data set shown below, complete parts (a) through (d) below. x] 20 30 40 50 600 y 100 95 89 83 70 (a) Use technology t
Answer #1

S.No (x-x)? 400.0000 100.0000 0.0000 100.0000 400.0000 1000.0000 SSX (y-7) 158.7600 57.7600 2.5600 19.3600 302.7600 541.2000

\\

a)

bo =116.20

b1=-0.72

b)

SSE =Syy-(Sxy)2/Sxx= 22.800000
error Variance σ2 = s2 =SSE/(n-2) = 7.6000
std error σ              = se =√s2 = 2.7568

c)

estimated standard error of slope   =se(β1) = s/√Sxx= 0.0872

d)

test stat t = β1/se(β1)= = -8.2590
p value: = 0.004